How can you describe the specific heat of a substance if it takes approximately 420 J of heat to raise the temperature of 7 g of olive oil by 10 °C?
1
Expert's answer
2020-05-27T13:34:20-0400

Solution.

"m=7g=0.007kg;"

"\Delta t=10^oC;"

"Q=420J;"

The specific heat of a substance is the amount of heat that must be provided to a substance weighing 1 kg to heat it by 1oC.

"c=\dfrac{Q}{m\Delta t};"

"c=\dfrac{420J}{0.007kg\sdot10^oC}=6000\dfrac{J}{kg^oC};"

Answer: "c=6000\dfrac{J}{kg^oC};"